### v4

# Observer Design for Networked Linear Systems with Fast and Slow Dynamics under Measurement Noise

Abstract
This paper addresses the emulation-based observer design for networked control systems (NCS) with linear plants that operate at two time scales in the presence of measurement noise. The system is formulated as a hybrid singularly perturbed dynamical system, enabling the systematic use of singular perturbation techniques to derive explicit bounds on the maximum allowable transmission intervals (MATI) for both fast and slow communication channels. Under the resulting conditions, the proposed observer guarantees that the estimation error satisfies a global exponential derivative-input-to-state stability (DISS)-like property, where the ultimate bound scales proportionally with the magnitudes of the measurement noise and the time derivative of the control input. The effectiveness of the approach is illustrated through a numerical example.
keywords:
Networked control systems, observer design, singular perturbed systems, hybrid dynamical systems
thanks: This work was supported by the Australian Research Council under the Discovery Project DP200101303, the France Australia collaboration project IRP-ARS CNRS and the ANR COMMITS ANR-23-CE25-0005, and by ANID through the FONDECYT Initiation grant 11250715. E-mail corresponding author: weixuanw@student.unimelb.edu.au.
1 Introduction
NCS, namely control systems whose loops are closed through packet-based communication networks, are popular across many domains. In some of these applications, the controlled systems feature multi–time-scale dynamics. For such systems, intuitively , signals that vary slowly can be sampled and transmitted at lower rates than those that change rapidly. However, most foundational works on NCS, e.g., carnevale_stability; yuksel2013stochastic, ignore the presence of multiple time scales and instead assume a uniform transmission rate, which lead unnecessarily high transmission rates for slow-varying signals and may impose avoidable communication burdens. Recent studies e.g., Romain_ETC; SPNCS; Single_channel_NCS_CDC; Weixuan_Journal, address the control of NCS with multiple time scales by applying singular perturbation techniques (nonlinear_systems_Khalil) to mitigate redundant transmissions of slow-varying signals and handle the inherent stiffness problem, leading to so called singularly perturbed NCS (SPNCS).
In addition to the challenges introduced by multiple time scales, many practical systems also involve internal states that cannot be directly measured. For example, in battery management, the state-of-charge and state-of-health are not directly measured and must instead be estimated. In such scenarios, observers are indispensable tools for estimating the unmeasurable states. Moreover, electrochemical, thermal and aging dynamics of battery evolve on multiple time scales, which naturally lead to the use of multi-time-scale observers (zou2016multi). Extensive research has been conducted on the multi-time-scale observer design for both continuous-time and sampled-data singularly perturbed systems, e.g., lin2010composite; yang2016observer; ferrante2021observer, while several studies also investigated observers for NCS, e.g., postoyan2011framework; wang2017observer. Despite these advances, the development of multi–time–scale observers for SPNCS that explicitly account for measurement noise, which frequently occurs in practice, remains unexplored.
In this work, we investigate the multi-time-scale observer design for SPNCS with linear plants, where the plant input, and both fast- and slow-varying measurements, are transmitted over a packet-based network subject measurement noise. Despite the linearity of the plant and observer, the problem remains nontrivial due to the potentially time-varying and nonlinear scheduling protocol, the hybrid nature of the closed-loop system, and network imperfections such as non-equidistant sampling.
Following an emulation-based approach, a Luenberger observer is first designed in continuous time while enhancing robustness against network-induced errors under the adopted uniformly globally exponentially stable (UGES) scheduling protocols (dragan_stability), with the resulting conditions expressed as linear matrix inequalities (LMI). This observer is then implemented over the network and the estimates on the maximum allowable transmission intervals (MATI) of slow and fast signals are provided, respectively. Under these conditions, the estimation error satisfies a global exponentially DISS-like property, in the sense of DISS. In particular, ultimate bound on the estimation error is proportional to the magnitudes of measurement noise and the time derivative of the plant input.
This work provides two main contributions. First, it introduces a systematic observer design framework for SPNCS with linear plants via LMI. In particular, the proposed observer explicitly incorporates key network-induced effects, including sampling and scheduling protocols. By doing so, it enables the derivation of observer gains that make the observer robust against network-induced errors and thereby maximize the MATI estimates for both slow- and fast-varying signals. This contribution provides a unified and computationally tractable method for observer synthesis in multi-time-scale NCS, where network constraints play a critical role. Second, the paper establishes a robustness analysis framework for observer design in SPNCS with linear plants subject to measurement noise and plant input. Because noisy measurements are sampled and transmitted over the network, the communication process inherently attenuates high-frequency noise. Following the idea in scheres2024robustifying, by analysing the received noise (i.e., the noise after it passes through the network), we remove the dependence of the ultimate bound on the time derivative of the measurement noise, leading to a more practical and implementable robustness guarantee.
Notation: Let $\mathbb{R}_{\geq 0}\coloneqq[0,\infty)$ and for any integer $n$, $\mathbb{Z}_{\geq n}\coloneqq\{n,n+1,\cdots\}$. For vectors $v_{i}\in\mathbb{R}^{n}$, $i\in\{1,2,\cdots,N\}$, we denote the vector $v_{1}^{\top}\;v_{2}^{\top}\;\cdots\;v_{N}^{\top}]^{\top}$ by $(v_{1},v_{2},\cdots,v_{N})$, and inner product by $\left$$\cdot$$\cdot\right$>. Given a vector $x\in\mathbb{R}^{n_{x}}$ and a non-empty closed set $\mathcal{A}\subseteq\mathbb{R}^{n_{x}}$, the distance from $x$ to $\mathcal{A}$ is denoted by $|x|_{\mathcal{A}}\coloneqq\min_{y\in\mathcal{A}}|x-y|$. Given a hybrid arc $u:\text{dom}(u)\rightarrow\mathcal{U}$ gosate12, define $\|u\|_{(t_{1},j_{1})}\coloneqq\max\Big\{\mathop{\mathrm{ess.sup}}\limits_{(t,j)\in\mathrm{dom}(u)\,\setminus\,\Gamma_{u}(u),\,(t,j)\preceq(t_{1},j_{1})}|u(t,j)|,$ $\mathop{\mathrm{sup}}\limits_{(t,j)\in\Gamma_{u}(u),\ (t,j)\preceq(t_{1},j_{1})}|u(t,j)|\Big\}$ where $\Gamma_{u}(u)$ denotes the set of all $(t,j)\in\text{dom}(u)$ such that $(t,j+1)\in\text{dom}(u)$. For a real symmetric matrix $P$, we denote its maximum and minimum eigenvalues by $\lambda_{\text{max}}(P)$ and $\lambda_{\text{min}}(P)$ respectively. The logic AND operator is denoted by $\wedge$.
2 Problem setting
This section outlines the structure of the plant $(\mathcal{P})$, observer $(\mathcal{O})$ and the communication network $(\mathcal{N})$ as depicted in Figure 1.
Figure 1: Block diagram of observer for SPNCS
2.1 Plant and observer
Consider the following linear singularly perturbed plant
$\mathcal{P}:\begin{cases}\begin{aligned} \dot{x}_{p}&=A_{11}x_{p}+A_{12}z_{p}+B_{1}u_{s}\\ \epsilon\dot{z}_{p}&=A_{21}x_{p}+A_{22}z_{p}+B_{2}u_{s}\\ \tilde{y}_{p_{s}}&=C_{1s}x_{p}+v_{1},\\ \tilde{y}_{p_{f}}&=C_{2s}x_{p}+C_{2f}z_{p}+v_{2},\end{aligned}\end{cases}$ (1)
where $0<\epsilon\ll 1$ is a known small parameter, $A_{22}$ is invertible and the pairs $(A_{11},C_{1s})$ and $(A_{22},C_{2f})$ are observable. Here, $x_{p}\in\mathbb{R}^{n_{x}}$ and $z_{p}\in\mathbb{R}^{n_{z}}$ denote the slow and fast states of the plant, respectively; $\tilde{y}_{p_{s}}\in\mathbb{R}^{n_{y_{s}}}$ and $\tilde{y}_{p_{f}}\in\mathbb{R}^{n_{y_{f}}}$ represent the slow and fast measured outputs; and $v_{1},v_{2}\in\mathcal{L}_{\infty}$ are measurement noise, where $n_{x_{p}},n_{z_{p}},n_{y_{s}},n_{y_{f}}\in\mathbb{Z}_{\geq 1}$. Moreover, $u_{s}$ is the “slow” input, and we assume it is differentiable with bounded derivative (in an $\mathcal{L}_{\infty}$ sense) when $\epsilon$ approaches zero (i.e., $u_{s}$ is absolutely continuous). This assumption is well justified in engineering practice, as most actuators cannot change instantaneously, i.e., they are rate-limited due to physical constraints such as inertia and friction, electrical restrictions on current and voltage, and safety considerations that prevent excessive accelerations.
For given $\epsilon$ in (1), the following observer is introduced:
$\mathcal{O}:\begin{cases}\begin{aligned} \dot{x}_{o}=&A_{11}x_{o}+A_{12}z_{o}+B_{1}\hat{u}_{s}\\ &+L_{1s}(\hat{\tilde{y}}_{p_{s}}-\hat{y}_{o_{s}})+L_{1f}(\hat{\tilde{y}}_{p_{f}}-\hat{y}_{o_{f}})\\ \epsilon\dot{z}_{o}=&A_{21}x_{o}+A_{22}z_{o}+B_{2}\hat{u}_{s}\\ &+L_{2s}(\hat{\tilde{y}}_{p_{s}}-\hat{y}_{o_{s}})+L_{2f}(\hat{\tilde{y}}_{p_{f}}-\hat{y}_{o_{f}})\\ y_{o_{s}}=&C_{1s}x_{o},\\ y_{o_{f}}=&C_{2s}x_{o}+C_{2f}z_{o},\end{aligned}\end{cases}$ (2)
where $x_{o}$ and $z_{o}$ are the estimated states. The variables $\hat{\tilde{y}}_{p_{s}}$ and $\hat{\tilde{y}}_{p_{f}}$ denote the most recently received noisy plant outputs transmitted over the network, while $\hat{y}_{o_{s}}$ and $\hat{y}_{o_{f}}$ are introduced as network-induced counterparts to ${y}_{o_{s}}$ and ${y}_{o_{f}}$, respectively, following the approach suggested in wang2017observer. For analysis purposes, we introduce auxiliary noise-free outputs $y_{p_{s}}=C_{1s}x_{p}$ and $y_{p_{f}}=C_{2s}x_{p}+C_{2f}z_{p}$. The observer gains $L_{1s}$, $L_{1f}$, $L_{2s}$ and $L_{2f}$ will be designed such that estimation error dynamics are robust to the network-induced error. In this work, we assume that zero-order-hold is applied to $\hat{\tilde{y}}_{p_{s}}$, $\hat{\tilde{y}}_{p_{f}}$ and $\hat{u}_{s}$ (i.e., $\dot{\hat{\tilde{y}}}_{p_{s}}=\dot{\hat{\tilde{y}}}_{p_{f}}=\dot{\hat{u}}_{s}=0$) between successive transmission instances.
2.2 Network
We define the measured network-induced errors as
$\tilde{e}_{p_{s}}\coloneqq\hat{\tilde{y}}_{p_{s}}-\tilde{y}_{p_{s}},\quad\tilde{e}_{p_{f}}\coloneqq\hat{\tilde{y}}_{p_{f}}-\tilde{y}_{p_{f}},\quad e_{u_{s}}\coloneqq\hat{u}_{s}-u_{s}.$ (3)
As shown in Figure 1, we consider the scenario in which $\tilde{y}_{p_{s}}$, $\tilde{y}_{p_{f}}$ share a common communication channel, while $u_{s}$ is transmitted over a dedicated channel. Each channel may consist of several network nodes, with each node exclusively handling either slow or fast signals. This configuration guarantees that slow and fast signals are never transmitted simultaneously. In particular, suppose there are $\ell$ slow nodes in the channel of plant output, then we can decompose $\tilde{e}_{p_{s}}=(\tilde{e}_{p_{s}}^{1},\cdots,\tilde{e}_{p_{s}}^{\ell})$ with each $k^{\text{th}}$ component having dimension $n_{k}$ corresponds to the $k^{\text{th}}$ slow node, while satisfying $\sum_{k=1}^{\ell}n_{k}=n_{y_{s}}$.
Let $\mathcal{T}\coloneqq\{t_{1},t_{2},\cdots\}$ represent the set of all transmission instants. Define the subset $\mathcal{T}^{s}\coloneqq\{t_{1}^{s},t_{2}^{s},\cdots\}$ as the instants when a slow node accesses the network, and let $\mathcal{T}^{f}\coloneqq\mathcal{T}\setminus\mathcal{T}^{s}=\{t_{1}^{f},t_{2}^{f},\cdots\}$ be the instants when a fast node transmits. For any $k\in\mathbb{Z}_{\geq 1}$, the transmission intervals satisfy
$\tau_{\text{miati}}^{s}\leq t_{k+1}^{s}-t_{k}^{s}\leq\tau_{\text{mati}}^{s},$ $\forall t_{k}^{s},t_{k+1}^{s}\in\mathcal{T}^{s},$ (4a)
$\tau_{\text{miati}}^{f}\leq t_{k+1}^{f}-t_{k}^{f}\leq\tau_{\text{mati}}^{f},$ $\forall t_{k}^{f},t_{k+1}^{f}\in\mathcal{T}^{f},$ (4b)
$\tau_{\text{miati}}^{f}\leq t_{k+1}-t_{k},$ $\forall t_{k},t_{k+1}\in\mathcal{T}.$ (4c)
Here, $0<\tau_{\text{miati}}^{f}\leq\tau_{\text{mati}}^{f}$ denote the minimum allowable transmission interval (MIATI) and MATI for consecutive fast transmissions, respectively, and $\tau_{\text{miati}}^{s}$ and $\tau_{\text{mati}}^{s}$ are defined similarly for slow transmissions. Since a slow transmission of $\tilde{y}_{p_{s}}$ may occur between two consecutive fast transmissions of $\tilde{y}_{p_{f}}$, the condition $\tau_{\text{miati}}^{f}\leq\tfrac{1}{2}\tau_{\text{mati}}^{f}$ is necessary for (4b) and (4c) to be valid.
Since the noisy measurements are sampled and transmitted, this process naturally filters out high-frequency noise. As a result, we only need to address the transmitted noise, rather than the original $v_{1}$ and $v_{2}$. For analytical purposes, we introduce the unmeasured noise-free variables $\hat{y}_{p_{s}}$, $\hat{y}_{p_{f}}$, $\hat{v}_{1}$ and $\hat{v}_{2}$, where $\hat{\star}$ denotes the most recently received $\star$ for all $\star\in\{y_{p_{s}},y_{p_{f}},v_{1},v_{2}\}$ like in scheres2024robustifying. Thus, we have $\hat{\tilde{y}}_{p_{s}}=\hat{y}_{p_{s}}+\hat{v}_{1}$ and $\hat{\tilde{y}}_{p_{f}}=\hat{y}_{p_{f}}+\hat{v}_{2}$. We define the ideal network-induced error
$e_{p_{s}}\coloneqq\hat{y}_{p_{s}}-y_{p_{s}},\quad e_{p_{f}}\coloneqq\hat{y}_{p_{f}}-y_{p_{f}},$
which leads to
$\tilde{e}_{p_{s}}=e_{p_{s}}+\hat{v}_{1}-v_{1},\quad\tilde{e}_{p_{f}}=e_{p_{f}}+\hat{v}_{2}-v_{2}.$
By adopting a similar approach introduced in postoyan2014tracking, we introduce artificial network-induced errors on the observer outputs, defined as
$e_{o_{s}}\coloneqq\hat{y}_{o_{s}}-y_{o_{s}},\quad e_{o_{f}}\coloneqq\hat{y}_{o_{f}}-y_{o_{f}}.$
The variables $\hat{y}_{o_{s}}$ and $\hat{y}_{o_{f}}$ are designed such that $e_{o_{s}}$ and $e_{o_{f}}$ evolve along the same vector field as $\tilde{e}_{p_{s}}$ and $\tilde{e}_{p_{f}}$ at transmission instants, respectively. To be explicit, we decompose the error and noise vectors as $e_{p_{s}}=(e_{p_{s}}^{1},\cdots,e_{p_{s}}^{\ell})$, $e_{o_{s}}=(e_{o_{s}}^{1},\cdots,e_{o_{s}}^{\ell})$ and $\hat{v}_{1}=(\hat{v}_{1}^{1},\cdots,\hat{v}_{1}^{\ell})$ with each $k^{\text{th}}$ component having dimension $n_{k}$ (i.e., same as $\tilde{e}_{p_{s}}$). When the slow node $j\in\{1,\cdots,\ell\}$ in $\tilde{e}_{p_{s}}$ accesses the channel at time $t_{k}^{s}\in\mathcal{T}^{s}$, the associated vectors are reset according to
[…]
for all $k\in\{1,\cdots,\ell\}$. Furthermore, we define the network-induced error on the ideal output estimation error as
$e_{y_{s}}\coloneqq e_{p_{s}}-e_{o_{s}},\quad e_{f}\coloneqq e_{p_{f}}-e_{o_{f}},$ (6)
so that $\hat{\tilde{y}}_{p_{s}}-\hat{y}_{o_{s}}=y_{p_{s}}-y_{o_{s}}+e_{y_{s}}+\hat{v}_{1}$ and $\hat{\tilde{y}}_{p_{f}}-\hat{y}_{o_{f}}=y_{p_{f}}-y_{o_{f}}+e_{f}+\hat{v}_{2}$.
At each transmission instance $t_{k}^{s}\in\mathcal{T}^{s}$ of a slow node, the network-induced errors are updated as follows: $\tilde{e}_{p_{s}}^{+}=h_{\tilde{p}s}(k,\tilde{e}_{p_{s}})$, $e_{us}^{+}=h_{us}(k,e_{us})$, $e_{y_{s}}^{+}=h_{y_{s}}(k,e_{y_{s}},e_{p_{s}})$, $\tilde{e}_{p_{f}}^{+}=\tilde{e}_{p_{f}}$, $e_{o_{f}}^{+}=e_{o_{f}}$, $e_{f}^{+}=e_{f}$. In addition, the variable $\hat{v}_{1}$ is updated via $\hat{v}_{1}^{+}=h_{v1}(k,\hat{v}_{1},v_{1},\tilde{e}_{p_{s}})$. Here, $h_{\tilde{p}s}$ and $h_{us}$ model the scheduling protocol that manage the slow updates of $\tilde{y}_{p_{s}}$ and $u_{s}$ as in dragan_stability, respectively, while $h_{y_{s}}$ and $h_{v1}$ are defined based on the update rule (5) and definition of $e_{y_{s}}$. Similarly, for each transmission time $t_{k}^{f}\in\mathcal{T}^{f}$ of a fast node, the updates are given by: $\tilde{e}_{p_{s}}^{+}=\tilde{e}_{p_{s}}$, $e_{us}^{+}=e_{us}$, $e_{y_{s}}^{+}=e_{y_{s}}$, $\tilde{e}_{p_{f}}^{+}=h_{\tilde{p}f}(\kappa_{f},\tilde{e}_{p_{f}})$, $e_{f}^{+}=h_{f}(\kappa_{f},e_{f},\tilde{e}_{p_{f}})$, and $\hat{v}_{2}^{+}=h_{v_{2}}(\kappa_{f},\hat{v}_{2},v_{2},\tilde{e}_{f})$.
3 A hybrid model for the SPNCS
$\begin{aligned} f_{\delta_{x}}(\delta_{x},e_{y_{s}},e_{u_{s}},\delta_{y},e_{f},\hat{v})\coloneqq&(A_{11}-L_{1s}C_{1s}-L_{1f}C_{2s})\delta_{x}+(A_{12}-L_{1f}C_{2f})(\delta_{y}+\overline{H}(\delta_{x},e_{y_{s}},e_{u_{s}},\hat{v}))\\ &+B_{1}e_{u_{s}}+L_{1s}(e_{y_{s}}+\hat{v}_{1})+L_{1f}(e_{f}+\hat{v}_{2})\end{aligned}$ (7)
[…]
$F_{\xi_{s}}(\xi_{s},\delta_{y},e_{f},\hat{v},\dot{u}_{s})\coloneqq\big(f_{\delta_{x}}(\delta_{x},e_{y_{s}},e_{u_{s}},\delta_{y},e_{f},\hat{v}),f_{e_{y_{s}}}(\delta_{x},e_{y_{s}},e_{u_{s}},\delta_{y},e_{f},\hat{v}),-\dot{u}_{s},1,0,0\big)$
$g_{\delta_{y}}(\delta_{x},e_{y_{s}},e_{u_{s}},\delta_{y},e_{f},\hat{v},\epsilon)\coloneqq(A_{22}-L_{2f}C_{2f})\delta_{y}+L_{2f}e_{f}-\epsilon\tfrac{\partial\overline{H}}{\partial\xi_{s}}F_{\xi_{s}}(\xi_{s},\delta_{y},e_{f},\hat{v},\dot{u}_{s})$
$g_{e_{f}}(\delta_{x},e_{y_{s}},e_{u_{s}},\delta_{y},e_{f},\hat{v},\epsilon)=\epsilon C_{2s}f_{\delta_{x}}(\delta_{x},e_{y_{s}},e_{u_{s}},\delta_{y},e_{f},\hat{v})+C_{2f}\big((A_{22}-L_{2f}C_{2f})\delta_{y}+L_{2f}e_{f}\big)$
$F_{\xi_{f}}(\xi_{s},\xi_{f},\hat{v},\dot{u}_{s},\epsilon)\coloneqq\big(g_{\delta_{y}}(\delta_{x},e_{y_{s}},e_{u_{s}},\delta_{y},e_{f},\hat{v},\epsilon),g_{e_{f}}(\delta_{x},e_{y_{s}},e_{u_{s}},\delta_{y},e_{f},\hat{v},\epsilon),1,0,0\big)$
$G_{s,1}(\xi_{s},\zeta_{s},v_{1})\coloneqq\big(\delta_{x},h_{ys}(\kappa_{s},e_{y_{s}},\tilde{e}_{p_{s}}),h_{us}(\kappa_{s},e_{u_{s}}),0,\kappa_{s}+1,h_{v_{1}}(\kappa_{s},\hat{v}_{1},v_{1},\tilde{e}_{p_{s}}),x_{p},h_{\tilde{p}s}(\kappa_{s},\tilde{e}_{p_{s}})\big)$
$G_{s,2}(\xi,v_{1})\coloneqq\big(h_{y}(\kappa_{s},e_{y_{s}},e_{u_{s}},\tilde{e}_{p_{s}},\delta_{y},\hat{v}_{1},v_{1}),e_{f},\tau_{f},\kappa_{f},\hat{v}_{2},z_{p},\tilde{e}_{p_{f}}\big)$
$\begin{aligned} G_{f}(\xi,v_{2})\coloneqq\big(&\delta_{x},e_{y_{s}},e_{u_{s}},\tau_{s},\kappa_{s},\hat{v}_{1},x_{p},\tilde{e}_{p_{s}},\delta_{y},h_{f}(\kappa_{f},e_{f},\tilde{e}_{p_{f}}),0,\kappa_{f}+1,h_{v_{2}}(\kappa_{f},\hat{v}_{2},v_{2},\tilde{e}_{f}),z_{p},h_{\tilde{p}f}(\kappa_{f},\tilde{e}_{p_{f}})\big)\end{aligned}$
We extend the standard singular perturbation technique (nonlinear_systems_Khalil, Section 11.5) to hybrid systems. Specifically, we begin by determining the quasi-steady state followed by a coordinate transformation. We then investigate the stability of the SPNCS by analyzing both the boundary-layer (fast) and the reduced (slow) system.
3.1 Change of coordinates
Define the estimation errors on the slow and fast states as $\delta_{x}\coloneqq x_{o}-x_{p}$ and $\delta_{z}\coloneqq z_{o}-z_{p}$, respectively. Let
$v\coloneqq(v_{1},v_{2}),\quad\hat{v}\coloneqq(\hat{v}_{1},\hat{v}_{2}).$ (8)
Then, we have
[…]
Let
$A_{11}^{f}\coloneqq A_{22}-L_{2f}C_{2f}.$ (9)
[…]
$\overline{H}(\delta_{x},e_{u_{s}},e_{y_{s}},\hat{v})\coloneqq-{A_{11}^{f}}^{-1}\big[(A_{21}-L_{2s}C_{1s}-L_{2f}C_{2s})\delta_{x}\\ +L_{2s}(e_{y_{s}}+\hat{v}_{1})+L_{2f}\hat{v}_{2}+B_{2}e_{u_{s}}\big]$
[…]
$\delta_{y}^{+}=$ $\delta_{z}^{+}-\overline{H}(\delta_{x}^{+},e_{u_{s}}^{+},e_{y_{s}}^{+},\hat{v}^{+})$ (10)
$=$ $\delta_{y}+\overline{H}(\delta_{x},e_{y_{s}},e_{u_{s}},\hat{v})-\overline{H}(\delta_{x}^{+},e_{y_{s}}^{+},e_{u_{s}}^{+},\hat{v}^{+})$
$\eqqcolon$ $h_{y}(\kappa_{s},e_{y_{s}},e_{u_{s}},\tilde{e}_{p_{s}},\delta_{y},\hat{v}_{1},v_{1}).$
where $\overline{H}(\delta_{x}^{+},e_{u_{s}}^{+},e_{y_{s}}^{+},\hat{v}^{+})=\overline{H}\big(\delta_{x},h_{us}(\kappa_{s},e_{u_{s}}),h_{ys}(\kappa_{s},e_{y_{s}},$ $\tilde{e}_{p_{s}}),(h_{v_{1}}(\kappa_{s},\hat{v}_{1},v_{1},\tilde{e}_{p_{s}}),\hat{v}_{2})\big)$.
3.2 Full order system
To formulate the SPNCS as a hybrid model, we define the full state vector
$\xi\coloneqq(\xi_{s},\zeta_{s},\xi_{f},\zeta_{f})\in\mathbb{X},$
where
$\xi_{s}$ $\coloneqq(\delta_{x},e_{y_{s}},e_{u_{s}},\tau_{s},\kappa_{s},\hat{v}_{1}),\quad$ $\zeta_{s}\coloneqq(x_{p},\tilde{e}_{p_{s}}),$ (11)
$\xi_{f}$ $\coloneqq(\delta_{y},e_{f},\tau_{f},\kappa_{f},\hat{v}_{2}),\quad$ $\zeta_{f}\coloneqq(z_{p},\tilde{e}_{p_{f}}),$
with $e_{y_{s}}$ and $e_{f}$ defined in (6), $e_{u_{s}}$, $\tilde{e}_{p_{s}}$, $\tilde{e}_{p_{f}}$ defined in (3)
Here, $\tau_{s},\tau_{f}\in\mathbb{R}_{\geq 0}$ denote clocks that measure the elapsed time since the most recent slow and fast transmission, respectively, and $\kappa_{s},\kappa_{f}\in\mathbb{Z}_{\geq 0}$ are counters for the number of slow and fast transmissions. The state space is defined as $\mathbb{X}\coloneqq\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{y_{s}}}\times\mathbb{R}^{n_{u_{s}}}\times\mathbb{R}_{\geq 0}\times\mathbb{Z}_{\geq 0}\times\mathbb{R}^{n_{y_{s}}}\times\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{y_{s}}}\times\mathbb{R}^{n_{z}}\times\mathbb{R}^{n_{y_{f}}}\times\mathbb{R}_{\geq 0}\times\mathbb{Z}_{\geq 0}\times\mathbb{R}^{n_{y_{f}}}\times\mathbb{R}^{n_{z}}\times\mathbb{R}^{n_{y_{f}}}$, and the SPNCS is modeled by the hybrid system (gosate12) with hybrid input $u_{s}$ and $v$, where $v$ is defined in (8).
$\mathcal{H}:\left\{\begin{aligned} \dot{\xi}&=F(\xi,\hat{v},u_{s},\dot{u}_{s},\dot{v},\epsilon),\ \xi\in\mathcal{C}^{\epsilon},\\ {\xi}^{+}&\in G(\xi,v),\ \xi\in\mathcal{D}_{s}^{\epsilon}\cup\mathcal{D}_{f}^{\epsilon},\end{aligned}\right$. (12)
where $F(\xi,\hat{v},u_{s},\dot{u}_{s},\dot{v},\epsilon)=\big(F_{\xi_{s}}(\xi_{s},\delta_{y},e_{f},\hat{v},\dot{u}_{s}),\allowbreak F_{\zeta_{s}}(x_{p},\allowbreak z_{p},u_{s},\dot{v}_{1}),\tfrac{1}{\epsilon}F_{\xi_{f}}(\xi,\hat{v},\dot{u}_{s},\epsilon),\tfrac{1}{\epsilon}F_{\zeta_{f}}(x_{p},z_{p},u_{s},\dot{v}_{2},\epsilon)\big)$, and $F_{\xi_{s}}$ and $F_{\xi_{f}}$ are defined in (7). Moreover, $F_{\zeta_{s}}(x_{p},z_{p},u_{s},\dot{v}_{1})=(\dot{x}_{p},-C_{1s}\dot{x}_{p}-\dot{v}_{1})$ and $F_{\zeta_{f}}(x_{p},z_{p},u_{s},\dot{v}_{2},\epsilon)=(\dot{z}_{p},-\epsilon C_{2s}\dot{x}_{p}-C_{2f}\dot{z}_{p}-\epsilon\dot{v}_{2})$. The jump map $G$ is given by
$G(\xi)\coloneqq\left\{\begin{aligned} &G_{s}(\xi,v_{1}),\;\xi\in\mathcal{D}_{s}^{\epsilon}\setminus\mathcal{D}_{f}^{\epsilon},\\ &G_{f}(\xi,v_{2}),\;\xi\in\mathcal{D}_{f}^{\epsilon}\setminus\mathcal{D}_{s}^{\epsilon},\\ &\{G_{s}(\xi,v_{1}),G_{f}(\xi,v_{2})\},\;\xi\in\mathcal{D}_{s}^{\epsilon}\cap\mathcal{D}_{f}^{\epsilon},\end{aligned}\right$. (13)
where $G_{s}(\xi,v_{1})=(G_{s,1}(\xi_{s},\zeta_{s},v_{1}),G_{s,2}(\xi,v_{2}))$, and $G_{s,1}$, $G_{s,2}$ and $G_{f}$ are defined in (7). The jump maps $G_{s}$ and $G_{f}$ correspond to the transmissions of slow and fast signals, respectively. Let $\tau_{\text{mati}}^{f}=\epsilon T^{*}$ with $T^{*}\in\mathbb{R}_{>0}$ independent of $\epsilon$ and define $\tau_{\text{miati}}^{f}=a\tau_{\text{mati}}^{f}$ for some $a\in(0,\tfrac{1}{2}]$, satisfying the inequality $\tau_{\text{miati}}^{f}\leq\tfrac{1}{2}\tau_{\text{mati}}^{f}$. Since $\epsilon>0$, the condition $\epsilon\tau_{f}\in[\tau_{\text{miati}}^{f},\tau_{\text{mati}}^{f}]$ is equivalent to $\tau_{f}\in[aT^{*},T^{*}]$. Consequently, the jump and flow sets in (12) are defined as
$\mathcal{D}_{s}^{\epsilon}$ $\coloneqq\!\{\xi\in\mathbb{X}\;|\;\tau_{s}\in[\tau_{\text{miati}}^{s},\tau_{\text{mati}}^{s}]\wedge\tau_{f}\in[aT^{*},(1-a)T^{*}]\},$
$\mathcal{D}_{f}^{\epsilon}$ $\coloneqq\!\{\xi\in\mathbb{X}\;|\;\tau_{s}\in[\epsilon aT^{*},\tau_{\text{mati}}^{s}-\epsilon aT^{*}]\wedge\tau_{f}\in[aT^{*},T^{*}]\}.$
This construction of $G$ guarantees the outer semi-continuity and the local boundness of the jump map, which is one of the hybrid basic conditions (gosate12). The flow set is given by $\mathcal{C}^{\epsilon}\coloneqq\mathcal{D}_{s}^{\epsilon}\cup\mathcal{D}_{f}^{\epsilon}\cup\mathcal{C}_{a}^{\epsilon}\cup\mathcal{C}_{b}^{\epsilon}$, with $\mathcal{C}_{a}^{\epsilon}\coloneqq\{\xi\in\mathbb{X}\ |\ \tau_{s}\in[0,\epsilon aT^{*}]\wedge\epsilon\tau_{f}\in[0,\tau_{s}+\epsilon T^{*}-\epsilon aT^{*}]\}$ and $\mathcal{C}_{b}^{\epsilon}\coloneqq\{\xi\in\mathbb{X}\;|\;\tau_{s}\in[\epsilon aT^{*},\epsilon\tau_{f}+\tau_{\text{mati}}^{s}-\epsilon aT^{*}]\wedge\epsilon\tau_{f}\in[0,\epsilon aT^{*}]\}$.
Remark 1
The states $\zeta_{s}$ and $\zeta_{f}$ do not affect the dynamics of estimation errors during flow; they appear only in the jump map of $e_{y_{s}}$, $e_{f}$ and $\hat{v}$. If a static protocol such as round-robin is implemented, $\zeta_{s}$ and $\zeta_{f}$ decouple from the estimation errors and their dynamics need not be included in $\mathcal{H}$.
3.3 Boundary-layer system and reduced system
We first derive the boundary-layer system. Define the fast time scale as $\sigma\coloneqq\tfrac{t}{\epsilon}$, so that $\tfrac{\partial}{\partial\sigma}=\epsilon\tfrac{\partial}{\partial t}$. By setting $\epsilon=0$ in system (12), the corresponding flow set $\mathcal{C}^{0}$ is given by
$\mathcal{C}^{0}\coloneqq\{\xi\in\mathbb{X}\ |\ \tau_{s}\in[0,\tau_{\text{mati}}^{s}]\wedge\tau_{f}\in[0,T^{*}]\},$
with $\mathcal{D}_{s}^{0}$, $\mathcal{D}_{f}^{0}$ derived accordingly. In the fast subsystem, the slow dynamics are effectively frozen, and the condition $\tau_{s}\in[0,\tau_{\text{mati}}^{s}]$ is always satisfied. Thus, the evolution of the boundary-layer system $\mathcal{H}_{bl}$ is governed solely by $\tau_{f}$. Consequently, we write
$\mathcal{H}_{bl}\!:\!\left\{\begin{aligned} \tfrac{\partial\xi}{\partial\sigma}&=(0,0,F_{\xi_{f}}(\xi_{s},\xi_{f},\hat{v},\dot{u}_{s},0),\\ &\qquad\quad F_{\zeta_{f}}(x_{p},z_{p},u_{s},0)),\quad\xi\!\in\mathcal{C}_{bl}^{0},\\ {\xi}^{+}&=G_{f}(\xi,v_{2}),\qquad\qquad\qquad\,\xi\!\in\mathcal{D}_{f}^{0},\end{aligned}\right$. (14)
where $\mathcal{C}_{bl}^{0}\coloneqq\{\xi\in\mathbb{X}\ |\ \tau_{f}\in[0,T^{*}]\}$, $\mathcal{D}_{f}^{0}\coloneqq\{\xi\in\mathbb{X}\ |\ \tau_{f}\in[aT^{*},T^{*}]\}$, $F_{\xi_{f}}$ and $G_{f}$ are defined in (7), and $F_{\zeta_{f}}$ is defined under (12). For clarity in stating our assumptions, we explicitly state the dynamic component of the boundary-layer system as
$(\tfrac{\partial\delta_{y}}{\partial\sigma},\tfrac{\partial{e_{f}}}{\partial\sigma})=\begin{bmatrix}A_{11}^{f}&A_{12}^{f}\\ A_{21}^{f}&A_{22}^{f}\end{bmatrix}\begin{bmatrix}\delta_{y}\\ e_{f}\end{bmatrix}$,
where $A_{11}^{f}$ is defined in (9), $A_{12}^{f}\coloneqq L_{2f}$, $A_{21}^{f}\coloneqq C_{2f}A_{11}^{f}$ and $A_{22}^{f}\coloneqq C_{2f}L_{2f}$.
From the perspective of slow dynamics, the fast dynamics evolve infinitely fast and thus remain at their quasi-steady state. Consequently, the reduced system $\mathcal{H}_{r}$ is given by
$\mathcal{H}_{r}\!\coloneqq\!\left\{\begin{aligned} (\dot{\xi}_{s},\dot{\zeta}_{s})&=(F_{\xi_{s}}(\xi_{s},0,0,\hat{v},\dot{u}_{s}),\\ &\qquad F_{\zeta_{s}}(x_{p},\overline{z}_{p},u_{s},\dot{v}_{1})),\quad\xi\in\mathcal{C}_{r}^{0},\\ (\xi_{s}^{+},\zeta_{s}^{+})&=G_{s,1}(\xi_{s},\zeta_{s},v_{1}),\qquad\quad\xi\in\mathcal{D}_{s}^{0},\end{aligned}\right$. (15)
where $\overline{z}_{p}\coloneqq-A_{22}^{-1}(A_{21}x_{p}+B_{2}u_{s})$ is the quasi-steady-state of $z_{p}$, $\mathcal{C}_{r}^{0}\coloneqq\{\xi\in\mathbb{X}\ |\ \tau_{s}\in[0,\tau_{\text{mati}}^{s}]\}$ and $\mathcal{D}_{s}^{0}\coloneqq\{\xi\in\mathbb{X}\ |\ \tau_{s}\in[\tau_{\text{miati}}^{s},\tau_{\text{mati}}^{s}]\}$. Define the error vector and its derivative
$e_{s}\coloneqq(e_{y_{s}},e_{u_{s}})\text{, }\dot{e}_{s}=f_{es}(\delta_{x},e_{y_{s}},e_{u_{s}},\delta_{y},e_{f},\hat{v},\dot{u}_{s}).$ (16)
Furthermore, we denote its update rule at slow transmission as
$h_{s}(\kappa_{s},e_{s},\tilde{e}_{p_{s}})\coloneqq\big(h_{ys}(\kappa_{s},e_{y_{s}},\tilde{e}_{p_{s}}),h_{u_{s}}(\kappa_{s},e_{u_{s}})\big).$
Then, a portion of the dynamics of the reduced system is given by
$\left[\begin{smallmatrix}\dot{\delta}_{x}\\ \dot{e}_{s}\end{smallmatrix}\right]=\left[\begin{smallmatrix}A_{11}^{s}&A_{12}^{s}\\ A_{21}^{s}&A_{22}^{s}\end{smallmatrix}\right]\left[\begin{smallmatrix}\delta_{x}\\ e_{s}\end{smallmatrix}\right]+\left[\begin{smallmatrix}A_{13}^{s}\\ A_{23}^{s}\end{smallmatrix}\right]\left[\begin{smallmatrix}\hat{v}_{1}\\ \hat{v}_{2}\end{smallmatrix}\right]+(\mathbf{0}_{n_{x}\times 1},\left[\begin{smallmatrix}\mathbf{0}_{n_{y_{s}}}\times 1\\ -\dot{u}_{s}\end{smallmatrix}\right]),$
where $A_{11}^{s}\coloneqq A_{11}^{r}$, $A_{12}^{s}\coloneqq\left[\begin{smallmatrix}A_{12}^{r}&A_{13}^{r}\end{smallmatrix}\right]$, $A_{13}^{s}\coloneqq\left[\begin{smallmatrix}A_{14}^{r}&A_{15}^{r}\end{smallmatrix}\right]$, $A_{21}^{s}\coloneqq\left[\begin{smallmatrix}A_{21}^{r}&\mathbf{0}\end{smallmatrix}\right]^{\top}$, $A_{22}^{s}\coloneqq\left[\begin{smallmatrix}A_{22}^{r}&A_{23}^{r}\\ \mathbf{0}&\mathbf{0}\end{smallmatrix}\right]$, $A_{23}^{s}\coloneqq\left[\begin{smallmatrix}A_{24}^{r}&A_{25}^{r}\\ \mathbf{0}&\mathbf{0}\end{smallmatrix}\right]$, $A_{11}^{r}\coloneqq(A_{11}-L_{1s}C_{1s}-L_{1f}C_{2s})-D(A_{21}-L_{2s}C_{1s}-L_{2f}C_{2s})$, $A_{12}^{r}\coloneqq L_{1s}-DL_{2s}$, $A_{13}^{r}\coloneqq B_{1}-DB_{2}$, $A_{14}^{r}\coloneqq A_{12}^{r}$, $A_{15}^{r}\coloneqq L_{1f}-DL_{2f}$, $A_{2\star}^{r}\coloneqq C_{1s}A_{1\star}^{r}$ for all $\star\in\{1,\cdots,5\}$ and $D\coloneqq(A_{12}-L_{1f}C_{2f})(A_{22}-L_{2f}C_{2f})^{-1}$.
4 Stability analysis
We begin by assuming that the scheduling protocol $h_{\star}$ is UGES with the associate Lyapunov function, as suggested in (dragan_stability, Definition 7). Let $n_{e_{s}}=n_{y_{s}}+n_{u_{s}}$ and $n_{e_{f}}=n_{y_{f}}$, we propose the following assumption on our scheduling protocol.
Assumption 1
For all $\star\in\{s,f\}$, there exist a function $W_{\star}:\mathbb{Z}_{\geq 0}\times\mathbb{R}^{n_{e_{\star}}}\to\mathbb{R}_{\geq 0}$ that is locally Lipschitz in its second argument. Moreover, there exist constants $\underline{a}_{W_{\star}},\overline{a}_{W_{\star}}>0$ as well as a constant $\lambda_{\star}\in[0,1)$ such that, for all $\kappa_{\star}\in\mathbb{Z}_{\geq 0}$ and $e_{\star}\in\mathbb{R}^{n_{e_{\star}}}$, the following properties hold:
$\underline{a}_{W_{\star}}\left|{e_{\star}}\right|\leq{W_{\star}}(k_{\star},{e_{\star}})\leq\overline{a}_{W_{\star}}\left|{e_{\star}}\right|,$ (17)
${W_{\star}}(\kappa_{\star}+1,h_{\star}(\kappa_{\star},e_{\star},\tilde{e}_{p_{\star}}))\leq\lambda_{\star}{W_{\star}}(\kappa_{\star},{e_{\star}}).$ (18)
Moreover, there exists $M_{\star}\geq 0$, such that for all $\kappa_{\star}\in\mathbb{Z}_{\geq 0}$ and almost all ${e_{\star}}\in\mathbb{R}^{n_{e_{\star}}}$, $\left|\tfrac{\partial W_{\star}(\kappa_{\star},e_{\star})}{\partial e_{\star}}\right|\leq M_{\star}$.
By definition of $f_{es}$ in (16), there exist $A_{H_{s}}=M_{s}A_{21}^{s}$ and $L_{s}=M_{s}|A_{22}^{s}|\underline{a}_{W_{s}}^{-1}$, such that for all $\delta_{x}\in\mathbb{R}^{n_{x}},\kappa_{s}\in\mathbb{Z}_{\geq 0}$ and almost all ${e_{s}}\in\mathbb{R}^{n_{e_{s}}}$, the following inequality holds,
$\left<\tfrac{\partial{W_{s}}(\kappa_{s},{e_{s}})}{\partial{e_{s}}},f_{es}(\delta_{x},e_{ys},e_{us},0,0,\hat{v},\dot{u}_{s})\right>\leq\\ L_{s}{W_{s}}(\kappa_{s},e_{s})+|A_{H_{s}}\delta_{x}|+\sigma_{w_{s}}(|\hat{v}_{1}|,|\hat{v}_{2}|,|\dot{u}_{s}|),$ (19)
where
$\sigma_{w_{s}}(|\hat{v}_{1}|,|\hat{v}_{2}|,|\dot{u}_{s}|)\coloneqq a_{w_{s}}^{v_{1}}|\hat{v}_{1}|+a_{w_{s}}^{v_{2}}|\hat{v}_{2}|+M_{s}|\dot{u}_{s}|,$ (20)
$a_{w_{s}}^{v_{1}}=M_{s}\left|A_{24}^{r}\right|$ and $a_{w_{s}}^{v_{2}}=M_{s}\left|A_{25}^{r}\right|$. This inequality bounds the growth of network-induced error during the flow of the reduced system.
Similarly, by definition of $g_{ef}$ in (7), there exist $L_{f}=M_{f}|a_{22}^{f}|\underline{a}_{W_{f}}^{-1}$ and $A_{H_{f}}=M_{f}A_{21}^{f}$ such that for all $x\in\mathbb{R}^{n_{x}},\kappa_{f}\in\mathbb{Z}_{\geq 0}$, and almost all ${e_{f}}\in\mathbb{R}^{n_{y_{f}}}$, the following inequality holds:
$\left<\tfrac{\partial{W_{f}}(\kappa_{f},{e_{f}})}{\partial{e_{f}}},g_{e_{f}}(\delta_{x},e_{ys},e_{us},\delta_{y},e_{f},\hat{v},0)\right>\leq\\ L_{f}{W_{f}}(\kappa_{f},e_{f})+|A_{H_{f}}\delta_{y}|.$ (21)
The following two assumptions guide a design procedure aimed at minimizing the gain between the network-induced error and the estimation error from the observer. A smaller gain enhances robustness to network-induced errors, allowing for slower transmission without compromising performance. Another approach is to directly maximize the transmission interval, as demonstrated in our numerical example. Let $P_{s}$, $P_{f}$ be positive definite symmetric matrices, and $a_{\rho_{s}}$, $\gamma_{s}$, $\eta_{1}$, $a_{\rho_{f}}$, $\gamma_{f}>0$. We introduce the following design condition, which is instrumental in reducing conservatism in the resulting MATI.
Condition 1
The variables $L_{2f}$, $P_{f}$ and $a_{\rho_{f}}$ are chosen so that (22) holds with the minimum feasible value of $\gamma_{f}$.
$\left[\begin{smallmatrix}P_{f}A_{11}^{f}+A_{11}^{f\top}P_{f}+a_{\rho_{f}}I+A_{H_{f}}^{\top}A_{H_{f}}&P_{f}A_{12}^{f}\\ A_{12}^{f\top}P_{f}&a_{\rho_{f}}\overline{a}_{W_{f}}^{2}I-\gamma_{f}^{2}\underline{a}_{W_{f}}^{2}I\end{smallmatrix}\right]\leq 0.$ (22)
The LMI (22) is always feasible since the pair $(A_{22},C_{2f})$ is observable. In particular, one can select $L_{2f}$ such that $A_{11}^{f}$ is Hurwitz, then there always exist $P_{f}>0$ such that
$P_{f}A_{11}^{f}+A_{11}^{f\top}P_{f}+a_{\rho_{f}}I+A_{H_{f}}^{\top}A_{H_{f}}<0.$ (23)
Moreover, by (23), we have
$(P_{f}A_{11}^{f}+A_{11}^{f\top}P_{f}+a_{\rho_{f}}I+A_{H_{f}}^{\top}A_{H_{f}})^{-1}<0,$
which also implies
$A_{12}^{f\top}P_{f}(A_{11}^{f}+A_{11}^{f\top}P_{f}+a_{\rho_{f}}I+A_{H_{f}}^{\top}A_{H_{f}}P_{f})^{-1}P_{f}A_{12}^{f}\leq 0.$
Since we can always find sufficiently large $\gamma_{f}$ such that
$(a_{\rho_{f}}\overline{a}_{W_{f}}^{2}I-\gamma_{f}^{2}\underline{a}_{W_{f}}^{2}I)\\ -A_{12}^{f\top}P_{f}(A_{11}^{f}+A_{11}^{f\top}P_{f}+a_{\rho_{f}}I+A_{H_{f}}^{\top}A_{H_{f}}P_{f})^{-1}P_{f}A_{12}^{f}\leq 0,$
the LMI (22) is satisfied by Schur complement.
Condition 2
Let $L_{2f}$ comes from Assumption 22. Then, $L_{1s}$, $L_{1f}$, $L_{2s}$, $P_{s}$, $a_{\rho_{s}}$ and $\eta_{1}$ are chosen such that (24) holds with the minimum feasible value of $\gamma_{s}$.
$\left[\begin{smallmatrix}P_{s}A_{11}^{s}+A_{11}^{s\top}P_{s}+\eta_{1}P_{s}^{\top}P_{s}+a_{\rho_{s}}I+A_{H_{s}}^{\top}A_{H_{s}}&P_{s}A_{12}^{s}\\ A_{12}^{s\top}P_{s}&a_{\rho_{s}}\overline{a}_{W_{s}}^{2}I-\gamma_{s}^{2}\underline{a}_{W_{s}}^{2}I\end{smallmatrix}\right]\leq 0.$ (24)
The feasibility of (24) can be checked along the same line as (22). The proof of Theorem 1 will show that inequality (24) implies $\mathcal{L}_{2}$ stability of reduced system (15) from $W_{s}(\kappa_{s},e_{s})$ to $|A_{Hs}\delta_{x}|$ with gain $\gamma_{s}$, and input-to-state stability of (15) w.r.t. $\hat{v}$. Similarly, (22) guarantee the $\mathcal{L}_{2}$ stability of the boundary-layer system (14).
Before stating our main result, we introduce the function $T(\cdot,\cdot,\cdot)$ (carnevale_stability), which provides an upper bound of $\tau_{\text{mati}}^{s}$ and $T^{*}$. Let
$T(L,\gamma,\lambda)\coloneqq\begin{cases}\tfrac{1}{Lr_{1}}\tan^{-1}(r_{2}),\quad\gamma>L\\ \tfrac{1}{L}\tfrac{1-\lambda}{1+\lambda},\quad\qquad\;\;\;\;\;\;\;\gamma=L\\ \tfrac{1}{Lr_{1}}\tanh^{-1}(r_{2}),\;\;\;\;\gamma<L,\end{cases}$ (25)
where $r_{1}\coloneqq\sqrt{|\left(\tfrac{\gamma}{L}\right)^{2}-1|}$, $r_{2}\coloneqq\tfrac{r_{1}(1-\lambda)}{2\tfrac{\lambda}{1+\lambda}\big(\tfrac{\gamma}{L}-1\big)+1+\lambda}$. In the special case where $L=0$, this nonlinear mapping simplifies to $T(0,\gamma,\lambda)=\tfrac{1}{\gamma}\big(\tan^{-1}(\tfrac{1}{\lambda})-\tan^{-1}(\lambda)\big).$ Lastly, we define the following attractor
$\mathcal{E}\coloneqq\left\{\xi\in\mathbb{X}:\delta_{x}=0\wedge e_{s}=0\wedge\delta_{y}=0\wedge e_{f}=0\right\}$
for the upcoming stability analysis. At this point, all necessary prerequisites have been established. We now state our main result.
Theorem 1
Consider the system (12) satisfying Assumptions 1, and Condition 22 and 24. Let $L_{s}$ and $L_{f}$ come from (19) and (21), respectively, $\gamma_{f}$ and $\gamma_{s}$ come from Conditions 22 and 24, respectively. Moreover, let $\lambda_{s}$ and $\lambda_{f}$ come from Assumption 1, and $\lambda_{f}^{*}\in[\lambda_{f},1)$. Then, for any $\tau_{\text{miati}}^{s}\leq\tau_{\text{mati}}^{s}<T(L_{s},\gamma_{s},\lambda_{s})$ and $\tau_{\text{mati}}^{f}\leq\epsilon T(L_{f},\gamma_{f},\lambda_{f}^{*})$, the following holds:
There exist $k,a,\epsilon^{*}>0$ and $\gamma_{v_{1}},\gamma_{v_{2}},\gamma_{\dot{u}_{s}}\geq 0$ such that for all $\epsilon\in(0,\epsilon^{*}]$, any solution $\xi$ to $\mathcal{H}$ with hybrid inputs $u_{s}$, $v_{1}$ and $v_{2}$, where $u_{s}$ is absolutely continuous and $\dot{u}_{s},v_{1},v_{2}\in\mathcal{L}_{\infty}$, satisfies a DISS property, i.e.,
$|\xi(t,j)|_{\mathcal{E}}\leq k|\xi(0,0)|_{\mathcal{E}}e^{-a(t+j)}+\gamma_{v_{1}}\|v_{1}\|_{(t,j)}\\ +\gamma_{v_{2}}\|v_{2}\|_{(t,j)}+\gamma_{\dot{u}_{s}}\|\dot{u}_{s}\|_{(t,j)},$ (26)
for any $(t,j)\in\text{dom}(\xi)$.
Proof: See Appendix A.
Remark 2
It is important to note that our result, presented in Theorem 1, shows that the ultimate bound depends on the derivative of the input signal and the magnitude of the measurement noise. This result is a special case of the standard DISS introduced in DISS.
Remark 3
A fundamental tradeoff exists between the MATI estimates and the ultimate bounds (i.e., $\gamma_{v_{1}}$, $\gamma_{v_{2}}$, and $\gamma_{\dot{u}_{s}}$). In particular, achieving larger MATI estimates typically results in increased ultimate bounds. This reflects the fact that allowing sparser transmissions reduces communication demands but at the expense of degraded robustness, as the network-induced errors and disturbance effects accumulate more significantly between transmissions.
5 Illustrative example
In this section, we provide a numerical example to illustrate the procedure for deriving the MATI and the corresponding ultimate bounds. We also present a simulation that demonstrates how variations in the control input influence the error dynamics, thereby highlighting the practical implications of the theoretical results.
Consider plant (1) with $A_{11}=\left[\begin{smallmatrix}a_{1}&0\\ 0&-a_{5}\end{smallmatrix}\right]$, $A_{12}=\left[\begin{smallmatrix}a_{2}&0\\ 0&0\end{smallmatrix}\right]$, $A_{21}=\left[\begin{smallmatrix}0&0\\ a_{3}&0\end{smallmatrix}\right]$, $A_{22}=\left[\begin{smallmatrix}-a_{2}&0\\ -a_{2}&-a_{4}\end{smallmatrix}\right]$, $C_{1s}=\left[\begin{smallmatrix}c_{1}&c_{2}\end{smallmatrix}\right]$, $C_{2s}=\left[\begin{smallmatrix}-1&0\end{smallmatrix}\right]$, $C_{2f}=\left[\begin{smallmatrix}0&-1\end{smallmatrix}\right]$, where $a_{1}=10^{-3}$, $a_{2}=0.37$, $a_{3}=1.1$, $a_{4}=4.9$, $a_{5}=1.2$, $c_{1}=-0.03$ and $c_{2}=1.9$. To simplify the computation, we assume observer gains have the form $L_{1s}=\left[\begin{smallmatrix}n_{1}\\ 0\end{smallmatrix}\right]$, $L_{2s}=\left[\begin{smallmatrix}-n_{2}\\ -n_{2}\end{smallmatrix}\right]$, $L_{1f}=\left[\begin{smallmatrix}0\\ n_{3}\end{smallmatrix}\right]$, and $L_{2f}=\left[\begin{smallmatrix}0\\ 0\end{smallmatrix}\right]$.
Since $u_{s}$, $y_{p_{s}}$ and $y_{p_{f}}$ are scalars, the protocols are given by $h_{s}(\kappa_{s},e_{s})=0$ and $h_{f}(\kappa_{f},e_{f})=0$, which are UGES protocols. Define $W_{\star}(\kappa_{\star},e_{\star})\coloneqq|e_{\star}|$, $s\in\{s,f\}$. Then, Assumption 1 holds with $\underline{a}_{W_{\star}}(s)=\overline{a}_{W_{\star}}(s)=s$, $\lambda_{\star}=0$ and $M_{\star}=1$. Moreover, inequality (19) is satisfied with $L_{s}=1.85$ and $A_{H_{s}}=A_{21}^{s}$, and (21) holds with $L_{f}=0$ and $A_{H_{f}}=A_{21}^{f}$.
By solving the LMI (22) and (24) simultaneously while maximizing $\epsilon^{*}T^{*}$, we obtain the parameters in observer gains to be $n_{1}=0.02$, $n_{2}=0.01$ and $n_{3}=0.18$. Meanwhile, we have $P_{s}=\left[\begin{smallmatrix}0.14&-0.55\\ -0.55&2.8\end{smallmatrix}\right]$, $\gamma_{s}=3.4$, $a_{\rho_{s}}=0.14$, $P_{f}=\left[\begin{smallmatrix}3.5&0.15\\ 0.15&2.7\end{smallmatrix}\right]$ and $\gamma_{f}=1.6$, $a_{\rho_{f}}=2.5$. We select $\tau_{\text{mati}}^{s}=0.15\ s<T(L_{s},\gamma_{s},\lambda_{s})$. Then, with $\lambda_{s}^{*}=0.33$, $\eta_{s1}=0.1$, $\eta_{s2}=0.1$, $\eta_{s3}=0.5$ and $\eta_{s}=\eta_{s1}+\eta_{s2}+\eta_{s3}=0.7$, Lemma 1 is satisfied (see Section A for details). Furthermore, we pick $\lambda_{f}^{*}=0.456$, then $T^{*}=T(L_{f},\gamma_{f},\lambda_{f}^{*})=0.45\ \nicefrac{{s}}{{\epsilon}}$.
For the given $P_{f}$ and $W_{s}$, the minimum possible values for $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, $\lambda_{4}$ and $\lambda_{5}$ in (32) are given by $0.0025$, $0.11$, $0.01$, $0.23$ and $0.01$, respectively. Similarly, for the specified $W_{s}$, $P_{s}$, $W_{f}$ and $P_{f}$, it can be shown that $b_{1}=14$, $b_{2}=0.21$, $b_{3}=2.8$, $a_{\Delta_{1}}=0.036$, $a_{\Delta_{2}}=0$ and $a_{\Delta_{3}}=0.056$. By selecting $\tau_{\text{miati}}^{s}=0.149\ ms$, $\mu=0.6a_{s}\underline{a}_{U_{s}}$ and $\mu_{1}=0.4a_{s}\underline{a}_{U_{s}}$, we obtain $\epsilon^{*}=0.016$, thus the maximum $\tau_{\text{mati}}^{f}$ is 7 ms. Moreover, we have $\gamma_{v_{1}}=8.4$, $\gamma_{v_{2}}=8.0$ and $\gamma_{\dot{u}_{s}}=10.0$. In this example, we aim for the maximum MATI estimates, as a result, the gain ultimate bound is large, as we discussed in Remark 3.
To verify the bound (26), we conducted simulations for the example presented above using $\epsilon=0.016$, $\tau_{\text{mati}}^{s}=0.15~\text{s}$, and $\tau_{\text{mati}}^{f}=7~\text{ms}\leq\epsilon^{*}T^{*}$, with initial conditions $x_{p}(0)=(0.1,0.1)$, $z_{p}(0)=(-0.2,-0.2)$, and all other states initialized at the origin. The results are shown in Figures 2 and 3, where two scenarios are considered: (i) a nonzero control input with no measurement noise, and (ii) zero control input subject to measurement noise.
The first scenario, shown in Figure 2, corresponds to $v_{1}=v_{2}=0$, where four different control inputs, namely $u_{s}=t$, $u_{s}=0.5t+100$, $u_{s}=100$ and $u_{s}=0$, are applied. We observe that the input with the larger time derivative (i.e., $u_{s}=t$) results in a larger ultimate bound, even though the input with the smaller derivative (i.e., $u_{s}=0.5t+100$) has a greater magnitude. Furthermore, when $\dot{u}_{s}=0$ (as in the cases $u_{s}=100$ and $u_{s}=0$), $u_{s}$ has no influence on the observer performance. This behaviour is consistent with the bound in (26).
Figure 2: Evolution of $|\xi|_{\mathcal{E}}$ under different $u_{s}$
The second scenario, shown in Figure 3, sets $u_{s}=0$ and introduces measurement noise. Two noise signals, $v_{1}=0.04\sin(25t)$ and $v_{1}=0.02\sin(50t)$, are simulated while $v_{2}=0$. The results show that the measurement noise with larger amplitude yields a larger ultimate bound, even though the time derivatives of the two noise signals have equal magnitude. Overall, the system’s response to both control inputs and measurement noise is consistent with the bound in (26), as illustrated in Figures 2 and 3.
Figure 3: Evolution of $|\xi|_{\mathcal{E}}$ under different $v_{1}$
6 Conclusion
This paper investigates the observer design for two-time-scale singularly perturbed networked control systems with linear plants that operate over a single communication channel, subject to both plant inputs and measurement noise. Our analysis demonstrates that under appropriate conditions, the resulting estimation error converges exponentially to a neighborhood of the origin, where the ultimate bound depends explicitly on the magnitudes of the measurement noise and the time derivative of the control input.
Appendix A Proof of Theorem 1
We first analysis the behaviour of systems (14) and (15), then we analysis the trajectory of the overall SPNCS.
Reduced system:
We denote $W_{\star}(\kappa_{\star},e_{\star})$ by $W_{\star}$ for $\star\in\{s,f\}$. For the reduced system (15), we define Lyapunov function $V_{s}(\delta_{x})=\delta_{x}^{\top}P_{s}\delta_{x}$, where $P_{s}$ comes from Assumption 22. Then we have $\underline{a}_{Vs}|\delta_{x}|^{2}\leq V_{s}(\delta_{x})\leq\overline{a}_{Vs}|\delta_{x}|^{2}$, where $\underline{a}_{Vs}=\lambda_{\text{min}}(P_{s})$ and $\overline{a}_{Vs}=\lambda_{\text{max}}(P_{s})$. We can show
$\left<\tfrac{\partial V_{s}}{\partial\delta_{x}},f_{\delta_{x}}(\delta_{x},e_{ys},e_{us},0,0,\hat{v})\right>=\delta_{x}^{\top}(P_{s}A_{11}^{s}+A_{11}^{s\top})\delta_{x}\\ +\delta_{x}^{\top}P_{s}A_{12}^{s}e_{s}+e_{s}^{\top}A_{12}^{s\top}P_{s}\delta_{x}+2\delta_{x}^{\top}P_{s}A_{13}^{s}\hat{v}$, (27)
where
$2\delta_{x}^{\top}P_{s}A_{13}^{s}\hat{v}=2\delta_{x}^{\top}P_{s}\begin{bmatrix}A_{14}^{r}&A_{15}^{r}\end{bmatrix}\begin{bmatrix}\hat{v}_{1}\\ \hat{v}_{2}.\end{bmatrix}$
By completion of squares, for any $\eta_{1,1},\eta_{1,2}>0$, we have
$2\delta_{x}^{\top}P_{s}A_{14}^{r}\hat{v}_{1}\leq\eta_{1,1}\delta_{x}^{\top}P_{s}^{\top}P_{s}\delta_{x}+a_{V_{s}}^{v_{1}}|\hat{v}_{1}|^{2},$ (28)
$2\delta_{x}^{\top}P_{s}A_{15}^{r}\hat{v}_{2}\leq\eta_{1,2}\delta_{x}^{\top}P_{s}^{\top}P_{s}\delta_{x}+a_{V_{s}}^{v_{2}}|\hat{v}_{2}|^{2},$
where $a_{V_{s}}^{v_{1}}\coloneqq\tfrac{1}{\eta_{1,1}}\lambda_{\text{max}}(A_{14}^{r\top}A_{14}^{r})$ and $a_{V_{s}}^{v_{2}}\coloneqq\tfrac{1}{\eta_{1,2}}\lambda_{\text{max}}(A_{15}^{r\top}\allowbreak A_{15}^{r})$. Let $\eta_{1,1}+\eta_{1,2}=\eta_{1}$, where $\eta_{1}$ comes from (24). By (27) and (28), we can write
[…]
$=$ $\delta_{x}^{\top}(P_{s}A_{11}^{s}+A_{11}^{s\top}+\eta_{1}P_{s}^{\top}P_{s})\delta_{x}+\delta_{x}^{\top}P_{s}A_{12}^{s}e_{s}$
$+e_{s}^{\top}A_{12}^{s\top}P_{s}\delta_{x}+\sigma_{V_{s}}(|\hat{v}_{1}|,|\hat{v}_{2}|),$
where $\sigma_{V_{s}}(|\hat{v}_{1}|,|\hat{v}_{2}|)=a_{V_{s}}^{v_{1}}|\hat{v}_{1}|^{2}+a_{V_{s}}^{v_{2}}|\hat{v}_{2}|^{2}$. By (24), the following inequality hold:
[…]
$\leq$ $-a_{\rho_{s}}|\delta_{x}|^{2}-a_{\rho_{s}}W_{s}^{2}-|A_{H_{s}}\delta_{x}|^{2}+\gamma_{s}^{2}W_{s}^{2}+\sigma_{V_{s}}(|\hat{v}_{1}|,|\hat{v}_{2}|).$
Inequality (29) demonstrates that Assumption 24 establishes an $\mathcal{L}_{2}$ stability of $\mathcal{H}_{r}$ with gain $\gamma_{s}$ from $W_{s}$ to $|A_{H_{s}}\delta_{x}|$ carnevale_stability.
The function $T$ in (25) is designed such at for a function $\phi$ whose derivative given by $\dot{\phi}=-2L\phi-\gamma(\phi^{2}+1)$ carnevale_stability, and the time it takes to decrease from $1/\lambda$ to $\lambda$ is $T(L,\gamma,\lambda)$.
Lemma 1
Let $\tau_{\text{mati}}^{s}<T(L_{s},\gamma_{s},\lambda_{s})$ be given. There exist $\lambda_{s}^{*}\in(\lambda_{s},1)$ and $\eta_{s}>0$ such that the solution $\phi_{s}$ to
$\dot{\phi}_{s}=-2L_{s}\phi_{s}-\gamma_{s}[(1+\eta_{s})\phi_{s}^{2}+1],$ (30)
with $\phi_{s}(0)=1/\lambda_{s}^{*}$ satisfies $\phi_{s}(\tau_{\text{mati}}^{s})=\lambda_{s}^{*}$ and $\phi_{s}(s)\in[\lambda_{s}^{*},1/\lambda_{s}^{*}]$ for all $s\in[0,\tau_{\text{mati}}^{s}]$.
We define a Lyapunov function
$U_{s}(\xi_{s})=V_{s}(\delta_{x})+\gamma_{s}\phi_{s}(\tau_{s})W_{s}^{2}(\kappa_{s},e_{s}),$
which was introduced in carnevale_stability, where $\phi_{s}$ and $\lambda_{s}^{*}$ come from Lemma 1, and $\phi_{s}(0)=1/\lambda_{s}^{*}$. Then by Lemma 1 and (17), there exist $\underline{a}_{U_{s}}\coloneqq\min\{\underline{a}_{V_{s}},\gamma_{s}\lambda_{s}^{*}\underline{a}_{W_{s}}^{2}\}$ and $\overline{a}_{U_{s}}\coloneqq\max\{\overline{a}_{V_{s}},\gamma_{s}\tfrac{1}{\lambda_{s}^{*}}\overline{a}_{W_{s}}^{2}\}$, such that $\underline{a}_{U_{s}}|(\delta_{x},e_{s})|^{2}\leq U_{s}(\xi_{s})\leq\overline{a}_{U_{s}}|(\delta_{x},e_{s})|^{2}$. By (19) and (29), we have
$\left<\tfrac{\partial U_{s}}{\partial\xi_{s}},F_{\xi_{s}}(\xi_{s},0,0,\hat{v},\dot{u}_{s})\right$>
[…]
$+\tfrac{\partial\phi_{s}}{\partial\tau_{s}}\gamma_{s}W_{s}^{2}+2\gamma_{s}\phi_{s}W_{s}\tfrac{\partial W_{s}}{\partial e_{s}}f_{es}(\delta_{x},e_{ys},e_{us},0,0,\hat{v},\dot{u}_{s})$
$\leq$ $-a_{\rho_{s}}(|\delta_{x}|^{2}+W_{s}^{2}(\kappa_{s},e_{s}))-|A_{H_{s}}\delta_{x}|^{2}+\gamma_{s}^{2}W_{s}^{2}$
$+\sigma_{V_{s}}(|\hat{v}_{1}|,|\hat{v}_{2}|)+\tfrac{\partial\phi_{s}}{\partial\tau_{s}}\gamma_{s}W_{s}^{2}$
$+2\gamma_{s}\phi_{s}W_{s}[L_{s}W_{s}+|A_{H_{s}}\delta_{x}|+M_{s}\sigma_{w_{s}}(|\hat{v}_{1}|,|\hat{v}_{2}|,|\dot{u}_{s}|)].$
By the definition of $\sigma_{w_{s}}$ and completion of squares, for any $\eta_{s1},\eta_{s2},\eta_{s3}>0$, we have
$2\gamma_{s}\phi_{s}W_{s}M_{s}\sigma_{w_{s}}\leq(\eta_{s1}+\eta_{s2}+\eta_{s3})\gamma_{s}^{2}\phi_{s}^{2}W_{s}^{2}\\ +\left(\tfrac{{a_{w_{s}}^{v_{1}}}^{2}}{\eta_{s1}}|\hat{v}_{1}|^{2}+\tfrac{{a_{w_{s}}^{v_{2}}}^{2}}{\eta_{s2}}|\hat{v}_{2}|^{2}+\tfrac{M_{s}^{2}}{\eta_{s2}}|\dot{u}_{s}|^{2}\right).$
Let $\eta_{s1}+\eta_{s2}+\eta_{s3}=\eta_{s}$, where $\eta_{s}$ comes from Lemma 1. Moreover, we have
$2\gamma_{s}\phi_{s}W_{s}|A_{H_{s}}\delta_{x}|\leq\gamma_{s}^{2}\phi_{s}^{2}W_{s}^{2}+|A_{H_{s}}\delta_{x}|^{2}.$
Hence, by definition of $\dot{\phi}_{s}$ in (30), we have that for reduced system,
$\big<\!\tfrac{\partial U_{s}}{\partial\xi_{s}},F_{\xi_{s}}(\xi_{s},0,0,\hat{v},\dot{u}_{s})\!\big>\!\leq\!-a_{s}|(\delta_{x},e_{s})|^{2}+\sigma_{s}(|\hat{v}_{1}|,|\hat{v}_{2}|,|\dot{u}_{s}|),$ (31)
where $a_{s}\coloneqq a_{\rho_{s}}\text{min}\{1,\underline{a}_{W_{s}}^{2}\}$ and
$\sigma_{s}(|\hat{v}_{1}|,|\hat{v}_{2}|,|\dot{u}_{s}|)\coloneqq\left(\tfrac{{a_{w_{s}}^{v_{1}}}^{2}}{\eta_{s1}}|\hat{v}_{1}|^{2}+\tfrac{{a_{w_{s}}^{v_{2}}}^{2}}{\eta_{s2}}|\hat{v}_{2}|^{2}+\tfrac{M_{s}^{2}}{\eta_{s2}}|\dot{u}_{s}|^{2}\right)\\ +\sigma_{V_{s}}(|\hat{v}_{1}|,|\hat{v}_{2}|).$
By Lemma 1, at transmission times of the reduced system (i.e., $t\in\mathcal{T}^{s}$), we have
$U_{s}(\xi_{s}^{+})=$ $U_{s}\big((\delta_{x},h_{s}(\kappa_{s},e_{s}),0,\kappa_{s}+1)\big)$
$=$ $V_{s}(\delta_{x})+\gamma_{s}\phi_{s}(0)W_{s}^{2}(\kappa_{s}+1,h_{s}(\kappa_{s},e_{s}))$
$\leq$ $V_{s}(\delta_{x})+\tfrac{\gamma_{s}}{\lambda_{s}^{*}}\lambda_{s}W_{s}^{2}(\kappa_{s},e_{s})$
$\leq$ $U_{s}(\xi_{s}).$
Boundary-layer system:
For boundary-layer system (14), let $V_{f}(\delta_{y})=\delta_{y}^{\top}P_{f}\delta_{y}$, where $P_{f}$ comes from Assumption 22, then we have $\underline{a}_{Vf}|\delta_{y}|^{2}\leq V_{f}(\delta_{y})\leq\overline{a}_{Vf}|\delta_{y}|^{2}$, where $\underline{a}_{Vf}=\lambda_{\text{min}}(P_{f})$ and $\overline{a}_{Vf}=\lambda_{\text{max}}(P_{f})$. Similar to (29), we can show
$\left<\tfrac{\partial V_{f}}{\partial\delta_{y}},g_{\delta_{y}}(\delta_{x},e_{ys},e_{us},\delta_{y},e_{f},\hat{v},\dot{v}_{1},0)\right>\leq\\ -a_{\rho_{f}}|\delta_{y}|^{2}-a_{\rho_{f}}W_{f}^{2}(\kappa_{f},e_{f})-|A_{H_{f}}\delta_{y}|^{2}+\gamma_{f}^{2}W_{f}^{2}(\kappa_{f},e_{f}).$
We define Lyapunov function
$U_{f}(\xi_{f})=V_{f}(\delta_{y})+\gamma_{f}\phi_{f}(\tau_{f})W_{f}^{2}(\kappa_{f},e_{f}),$
where $\tfrac{\partial{\phi}_{f}}{\partial\sigma}=-2L_{f}\phi_{f}-\gamma_{f}(\phi_{f}^{2}+1)$, $\phi_{f}(0)=1/\lambda_{f}^{*}$ and $\lambda_{f}^{*}$ comes from Theorem 1. Given $T^{*}\leq T(L_{f},\gamma_{f},\lambda_{f}^{*})$, we have $\phi_{f}(s)\in[\lambda_{f}^{*},1/\lambda_{f}^{*}]$ for all $s\in[0,T^{*}]$. Then by (17), we can find $\underline{a}_{U_{f}}\coloneqq\min\{\underline{a}_{V_{f}},\gamma_{f}\lambda_{f}^{*}\underline{a}_{W_{f}}^{2}\}$ and $\overline{a}_{U_{f}}\coloneqq\max\{\overline{a}_{V_{f}},\gamma_{f}\tfrac{1}{\lambda_{f}^{*}}\overline{a}_{W_{f}}^{2}\}$ such that $\underline{a}_{U_{f}}|(\delta_{y},e_{f})|^{2}\leq U_{f}(\xi_{f})\leq\overline{a}_{U_{f}}|(\delta_{y},e_{f})|^{2}$. Along the same lines as for the reduced system, we can show
$\big<\tfrac{\partial U_{f}}{\partial\xi_{f}},F_{\xi_{f}}(\xi_{s},0,0,\hat{v},\dot{u}_{s})\big>\leq-a_{f}|(\delta_{y},e_{f})|^{2},$
where $a_{f}\coloneqq a_{\rho_{f}}\text{min}\{1,\underline{a}_{W_{f}}^{2}\}$. Moreover, at transmission time of $\mathcal{H}_{bl}$ (i.e., $t\in\mathcal{T}^{f}$), we can show $U_{f}(\xi_{f}^{+})\leq U_{f}(\xi_{f}$.
Overall SPNCS:
Now we focus on the full system $\mathcal{H}$. Remind that at transmission time of slow signals, we have $\delta_{y}$ updating according to (10). Then by (17) and definition of $V_{f}$, we can find $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4},\lambda_{5}\geq 0$ such that
$V_{f}(h_{y}(\kappa_{s},e_{ys},e_{us},\tilde{e}_{p_{s}},\delta_{y},\hat{v}_{1},v_{1}))-V_{f}(\delta_{y})$ (32)
$\leq$ $\lambda_{1}W_{s}^{2}(\kappa_{s},e_{s})+\lambda_{2}\sqrt{W_{s}^{2}(\kappa_{s},e_{s})V_{f}(\delta_{y})}$
[…]
Let $a_{\psi_{s}}\coloneqq\sqrt{\underline{a}_{U_{s}}}$, $a_{\psi_{f}}\coloneqq\sqrt{\underline{a}_{U_{f}}}$ and
$\mu\in(0,a_{s}a_{\psi_{s}}^{2}),\quad\mu_{1}\in(0,\mu).$ (33)
Let
$\tilde{\lambda}\in(e^{-\mu_{1}\tau_{\text{miati}}^{s}},1),$ (34)
where $\tau_{\text{miati}}^{s}$ comes from Theorem 1. We define
$d\coloneqq\left(\tfrac{-b+\sqrt{b^{2}-4ac}}{2a}\right)^{2},$ (35)
where $a=\max\{\tfrac{\lambda_{1}+\nicefrac{{\lambda_{5}}}{{2}}}{\gamma_{s}\lambda_{s}^{*}},\tfrac{\lambda_{4}}{2}\}$, $b=\tfrac{\lambda_{2}}{2}\max\{\tfrac{\lambda_{1}}{\gamma_{s}\lambda_{s}^{*}},1\}$ and $c=1-\tilde{\lambda}e^{\mu_{1}\tau_{\text{miati}}^{s}}$. Let
$a_{d}\coloneqq 1+a\cdot d+b\sqrt{d}.$ (36)
We define a composite Lyapunov function
$U(\xi_{s},\xi_{f})\coloneqq U_{s}(\xi_{s})+dU_{f}(\xi_{f}),$
and we can find $\underline{a}_{U},\overline{a}_{U}>0$ such that $\underline{a}_{U}|\xi|_{\mathcal{E}}^{2}\leq U(\xi)\leq\overline{a}_{U}|\xi|_{\mathcal{E}}^{2}$. We can show at a fast transmission, $U(\xi_{s}^{+},\xi_{f}^{+})\leq U(\xi_{s},\xi_{f})$. During flow, the time derivative of $U$ along the trajectories of $\mathcal{H}$ is given by
$\tfrac{\partial U(\xi_{s},\xi_{f})}{\partial t}$
$=$ $\tfrac{\partial U_{s}}{\partial\xi_{s}}F_{\xi_{s}}(\xi_{s},\delta_{y},e_{f},\hat{v},\dot{u}_{s})+\tfrac{d}{\epsilon}\tfrac{\partial U_{f}}{\partial\xi_{f}}F_{\xi_{f}}(\xi_{s},\xi_{f},\hat{v},\dot{u}_{s},\epsilon)$
$=$ $\tfrac{\partial U_{s}}{\partial\xi_{s}}F_{\xi_{s}}(\xi_{s},0,0,\hat{v},\dot{u}_{s})+\tfrac{d}{\epsilon}\tfrac{\partial U_{f}}{\partial\xi_{f}}F_{\xi_{f}}(\xi_{s},\xi_{f},\hat{v},\dot{u}_{s},0)$
$+\tfrac{\partial U_{s}}{\partial\xi_{s}}\left[F_{\xi_{s}}(\xi_{s},\delta_{y},e_{f},\hat{v},\dot{u}_{s})-F_{\xi_{s}}(\xi_{s},0,0,\hat{v},\dot{u}_{s})\right]$
$+\tfrac{d}{\epsilon}\tfrac{\partial U_{f}}{\partial\xi_{f}}\left[F_{\xi_{f}}(\xi_{s},\xi_{f},\hat{v},\dot{u}_{s},\epsilon)-F_{\xi_{f}}(\xi_{s},\xi_{f},\hat{v},\dot{u}_{s},0)\right].$
By definition of $U_{s}$, $F_{\xi_{s}}$ and Assumption 1 (i.e., $\left|\tfrac{\partial W_{s}}{\partial e_{s}}\right|\leq M_{s}$), there exists $b_{1}\geq 0$ such that
$\tfrac{\partial U_{s}}{\partial\xi_{s}}[F_{\xi_{s}}(\xi_{s},\delta_{y},e_{f},\hat{v},\dot{u}_{s})-F_{\xi_{s}}(\xi_{s},0,0,\hat{v},\dot{u}_{s})]\leq\\ b_{1}|(\delta_{x},e_{s})|\cdot|(\delta_{y},e_{f})|.$
Similarly, by definition of $U_{f}$, $F_{\xi_{f}}$, Assumption 1 and completion of squares, we can show there exist $b_{2},b_{3},a_{\Delta_{1}},a_{\Delta_{2}},\allowbreak a_{\Delta_{3}}\geq 0$ such that
$\tfrac{1}{\epsilon}\tfrac{\partial U_{f}}{\partial\xi_{f}}\left[F_{\xi_{f}}(\xi_{s},\xi_{f},\hat{v},\dot{u}_{s},\epsilon)-F_{\xi_{f}}(\xi_{s},\xi_{f},\hat{v},\dot{u}_{s},0)\right]\\ \leq b_{2}|(\delta_{x},e_{s})|\cdot|(\delta_{y},e_{f})|+b_{3}|(\delta_{y},e_{f})|^{2}+\Delta(|\hat{v}_{1}|,|\hat{v}_{2}|,|\dot{u}_{s}|),$
where $\Delta(|\hat{v}_{1}|,|\hat{v}_{2}|,|\dot{u}_{s}|)\coloneqq a_{\Delta_{1}}|\hat{v}_{1}|^{2}+a_{\Delta_{2}}|\hat{v}_{2}|^{2}+a_{\Delta_{3}}|\dot{u}_{s}|^{2}$. Then we have
$\tfrac{\partial U(\xi_{s},\xi_{f})}{\partial t}\leq-\left[\begin{smallmatrix}\sqrt{U_{s}(\xi_{s})}\\ \sqrt{U_{f}(\xi_{f})}\end{smallmatrix}\right]^{\top}\Lambda\left[\begin{smallmatrix}\sqrt{U_{s}(\xi_{s})}\\ \sqrt{U_{f}(\xi_{f})}\end{smallmatrix}\right]+\Gamma(|\hat{v}_{1}|,|\hat{v}_{2}|,|\dot{u}_{s}|),$
where
$\Lambda\coloneqq\left[\begin{smallmatrix}a_{s}a_{\psi_{s}}^{2}&-\tfrac{1}{2}(b_{1}+db_{2})a_{\psi_{s}}a_{\phi_{f}}\\ -\tfrac{1}{2}(b_{1}+db_{2})a_{\psi_{s}}a_{\psi_{f}}&d(\frac{a_{f}}{\epsilon}-b_{3})a_{\psi_{f}}^{2}\end{smallmatrix}\right],$
and
$\Gamma(|\hat{v}_{1}|,|\hat{v}_{2}|,|\dot{u}_{s}|)\coloneqq\sigma_{s}(|\hat{v}_{1}|,|\hat{v}_{2}|,|\dot{u}_{s}|)+d\cdot\Delta(|\hat{v}_{1}|,|\hat{v}_{2}|,|\dot{u}_{s}|).$
We define
$\epsilon^{*}=\left(\tfrac{1}{da_{f}a_{\psi_{f}}^{2}}\Big(\tfrac{(b_{1}+db_{2})^{2}a_{\psi_{s}}^{2}a_{\psi_{f}}^{2}}{4(a_{s}a_{\psi_{s}}^{2}-\mu)}+\mu d\Big)+\tfrac{b_{3}}{a_{f}}\right)^{-1},$
then, by definition of $d$ in (35) and $\mu$ in (33), we have $\Lambda\geq\mu\left[\begin{smallmatrix}1&0\\ 0&d\end{smallmatrix}\right]$, and by definition of $\mu_{1}$ in (33), we have
$\tfrac{\partial U(\xi_{s},\xi_{f})}{\partial t}\leq-\mu_{1}U(\xi_{s},\xi_{f}),\quad\forall U(\xi_{s},\xi_{f})\geq\tfrac{\Gamma(|\hat{v}_{1}|,|\hat{v}_{2}|,|\dot{u}_{s}|)}{\mu-\mu_{1}}.$
Let $t_{j_{k}^{s}}\in\mathcal{T}^{s}$ correspond to the transmission instance of the ${k^{s}}^{\text{th}}$ slow transmission, $j_{k}^{s}$ denotes the number of transmissions including the ${k^{s}}^{\text{th}}$ slow transmission at time $t_{j_{k}^{s}}$, and we denote $U(\xi_{s}(t,j),\xi_{f}(t,j))$ by $U(t,j)$. Then, by comparison principle and the fact that $U$ is non-increasing at fast transmission, we have that for all $t_{j_{k}^{s}},t_{j_{k+1}^{s}}\in\mathcal{T}^{s}$, if $U(t_{j_{k}^{s}},j_{k}^{s})\geq\tfrac{\Gamma(|\hat{v}_{1}|,|\hat{v}_{2}|,|\dot{u}_{s})}{\mu-\mu_{1}}$, then
$U(s,i)\leq U(t_{j_{k}^{s}},j_{k}^{s})\exp\big(-\mu_{1}(s-t_{j_{k}^{s}})\big),$
whenever $U(s,i)\geq\tfrac{\Gamma(|\hat{v}_{1}|,|\hat{v}_{2}|,|\dot{u}_{s})}{\mu-\mu_{1}}$, as well as $(t_{j_{k}^{s}},j_{k}^{s})\preceq(s,i)\preceq(t_{j_{k+1}^{s}},j_{k+1}^{s}-1)$.
At slow transmissions, we have
$U(\xi_{s}^{+},\xi_{f}^{+})\leq U_{s}(\xi_{s})+dV_{f}(h_{y}(\kappa_{s},e_{ys},e_{us},\tilde{e}_{p_{s}},\delta_{y},\hat{v}_{1},v_{1})).$
By adding and subtracting $V_{f}(\xi_{f})$, we have
$U(\xi_{s}^{+},\xi_{f}^{+})\leq U(\xi_{s},\xi_{f})+d\Big(\lambda_{1}W_{s}^{2}(\kappa_{s},e_{s})\\ +\lambda_{2}\sqrt{W_{s}^{2}(\kappa_{s},e_{s})V_{f}(\delta_{y})}+\lambda_{3}|\hat{v}_{1}|^{2}+\lambda_{4}\sqrt{V_{f}}|\hat{v}_{1}|+\lambda_{5}W_{s}|\hat{v}_{1}|\Big),$
and by completion of square, we can obtain
$U(\xi_{s}^{+},\xi_{f}^{+})\leq a_{d}U(\xi_{s},\xi_{f})+a_{v}|\hat{v}_{1}|^{2},$
where $a_{d}$ is defined in (36) and
$a_{v}\coloneqq d\lambda_{3}+\tfrac{\lambda_{4}+d\lambda_{5}}{2}.$
Then if $U(s,i)\geq\tfrac{\Gamma(|\hat{v}_{1}|,|\hat{v}_{2}|,|\dot{u}_{s}|)}{\mu-\mu_{1}}$ for all $(t_{j_{k}^{s}},j_{k}^{s})\preceq(s,i)\preceq(t_{j_{k+1}^{s}},j_{k+1}^{s}-1)$, then
$U(t^{s}_{k+1},j^{s}_{k+1})\leq\left(a_{d}U(t_{j_{k}^{s}},j_{k}^{s})+a_{v}|\hat{v}_{1}|^{2}\right)e^{-\mu_{1}\tau_{\text{miati}}^{s}}.$
By definition of $\tilde{\lambda}$ in (34), we have $a_{d}e^{-\mu_{1}\tau_{\text{miati}}^{s}}=\tilde{\lambda}$. Let $\lambda\in(\tilde{\lambda},1)$, then if $U(s,i)\geq\tilde{\Gamma}(|\hat{v}_{1}|,|\hat{v}_{2}|,|\dot{u}_{s}|)$ for all $(t_{j_{k}^{s}},j_{k}^{s})\preceq(s,i)\preceq(t_{j_{k+1}^{s}},j_{k+1}^{s}-1)$, we have
$U(t^{s}_{k+1},j^{s}_{k+1})\leq\lambda U(t_{j_{k}^{s}},j_{k}^{s}),$
where $\tilde{\Gamma}(|\hat{v}_{1}|,|\hat{v}_{2}|,|\dot{u}_{s}|)\coloneqq\max\{\tfrac{a_{v}|\hat{v}_{1}|^{2}}{\lambda-\tilde{\lambda}},\tfrac{\Gamma(|\hat{v}_{1}|,|\hat{v}_{2}|,|\dot{u}_{s}|)}{\mu-\mu_{1}}\}$.
Following a similar proof as in Weixuan_Journal and by comparison principle, we can show that
$|\xi(t,j)|_{\mathcal{E}}\leq\big(\tfrac{a_{d}\overline{a}_{U}}{\lambda\underline{a}_{U}}\big)^{\nicefrac{{1}}{{2}}}|\xi(0,0)|_{\mathcal{E}}\text{exp}\left(-\tfrac{\ln{(\nicefrac{{1}}{{\lambda}})}}{2\tau_{\text{mati}}^{s}}t\right)\\ +\gamma_{\hat{v}}\|\hat{v}\|_{(t,j)}+\gamma_{\dot{u}_{s}}\|\dot{u}_{s}\|_{(t,j)},$
where
$\gamma_{\hat{v}_{1}}$ $=\left(\tfrac{a_{d}}{\underline{a}_{U}}\max\left\{\tfrac{a_{v}}{\lambda-\tilde{\lambda}},\tfrac{1}{\mu-\mu_{1}}\big(\tfrac{{a_{ws}^{v_{1}}}^{2}}{\eta_{s1}}+a_{V_{s}}^{v_{1}}+da_{\Delta_{1}}\big)\right\}\right)^{\nicefrac{{1}}{{2}}}$
$\gamma_{\hat{v}_{2}}$ $=\left(\tfrac{a_{d}}{\underline{a}_{U}(\mu-\mu_{1})}\big(\tfrac{{a_{ws}^{v_{2}}}^{2}}{\eta_{s2}}+a_{V_{s}}^{v_{2}}+da_{\Delta_{2}}\big)\right)^{\nicefrac{{1}}{{2}}}$
$\gamma_{\dot{u}_{s}}$ $=\big(\tfrac{a_{d}}{\underline{a}_{U}}(\tfrac{M_{s}^{2}}{\eta_{s3}}+da_{\Delta_{3}})\big)^{\nicefrac{{1}}{{2}}}$
Moreover, we can always find $\gamma_{v_{\ell}}\geq 0$ such that $\gamma_{\hat{v}_{\ell}}\|\hat{v}_{\ell}\|_{(t,j)}\leq\gamma_{v_{\ell}}\|{v_{\ell}}\|_{(t,j)}$ for $\ell\in\{1,2\}$. For instance, $\gamma_{v_{\ell}}=\gamma_{\hat{v}_{\ell}}$ in sampled-data system. Finally, we prove Theorem 1 by using equation (A.25) in Weixuan_Journal.